What is the equation for the parabola with vertex (0, 0) and focus (5, 0)
Answer by lwsshak3(11628)

What is the equation for the parabola with vertex (0, 0) and focus (5, 0)
**
Based on given information, this is a parabola with an equation of the standard form:
(y-k)^2=4p(x-h), with (h,k) being the (x,y) coordinates of the vertex.
For given problem:
vertex: (0,0) given
p=5 (based on given focus)
parabola opens rightwards
Equation:
(y-0)^2=4p(x-0)
y^2=20x